# An Insight of Heart-Like Systems with Percolation

### v1

@@ Abstract
We study the signal percolation through heart-like biological system. Starting from an initial distribution of Waiting and Inactive cells with probabilities $p$ and $(1-p)$ respectively the signal propagation is observed in terms of Active cells. As the signal enters the system from one end, the number of arrival of Active sites at the other end is studied and analysis of the system behaviour is made by varying a few important parameters of the system like $p_{switch}$ (switching probability from Inactive to Waiting) and $p_{act}$ (switching probability from Waiting to active). In this connection, the non-regular heart rhythms are discussed. Fraction of percolating paths through the system is studied and showed a transition from 00 to $1$ near $p=p_{c}$. Some other important quantities like tortuosity and cluster distribution are discussed and corresponding exponents have been found.
@@ Introduction
Percolation is one of the simplest and fundamental models in statistical mechanics which is not always an exactly solvable model. In spite of its simple update rules, it can be applied to describe a wide variety of systems. Not only for the study of the dynamics of fluids through porous media [1] but it can as well be applied to study the electrical movement in neurons, fibrosis in organs, damage spreading in solids through fiber bundle model, disease spreading in a community, opinion formation in society or even for knowledge percolation [2, 3, 4, 5, 6, 7, 8, 9].
A very popular model of Percolation is the forest fire model (FFM) introduced in [10, 11] mainly as a 2D lattice. There are several variations of the model [11, 12]. In one, the fire can spread from top to bottom and from left to right can be regarded as a general “Directed Percolation” (DP) model. For DP, it is well understood that the bottom to top and right to left movements are prohibited [13, 14]. There are other models of percolation where all top to bottom, bottom to top, left to right and right to left movements are all allowed and can be termed as Isotropic Percolation (IP). However, the model where the percolation is allowed in the top to bottom direction and sideways can be termed as Semi Directed Percolation (SDP). This was introduced for the first time [15] in a random resistor-diode network in [16]. It has also been studied in square and triangular lattices [17]. However in both [16, 17], it was termed not as SDP but as Partially Directed Percolation.
@@
Several biological models have been used to study the heart system with the help of computation. Computational models in cardiology showed enormous scope in revealing diagnostic information related to cardiac disease. At the same time there is provisions that these models help the diagnosis to be cost-effective and to have low risk [5]. The cells can be activated by an external stimulus which initiates an Action Potential (AP). The cells after activation will go through a certain refractory period. The exact physiological processes are described in [22] in a detailed manner. In [6], the model of heart is studied with three states of cells: active, inactive and waiting in connection with the AP.
One of the simplified system to study a percolation problem is a $L\times L$ lattice, where there is a parameter called percolation probability $p$ with which a site could be open, i.e. can percolate “something”. This in turn indicates that a site is closed with $1-p$ probability, means unable to percolate. It is well known from different studies that below a critical percolation probability $p_{c}$ the system is unable to percolate, whereas above $p_{c}$ infinite percolation cluster can be formed. In connection to percolation there are a number of exponents which can be studied as well. Those exponents actually describes the system near criticality.
@@
We study a 2D square grid of $L\times L$ cells where initially the cells may be in two states, “Waiting” and “Inactive” as described in [6] also. A “Waiting” cell is an open cell which is ready to hold anything percolating through the system. An “Inactive” cell is however equivalent to a closed cell which is unable to percolate. We choose the system with an initial distribution of “Waiting” cells with probability $p$ and “Inactive” cells with probability $1-p$. A “Waiting” cell can be activated with an action potential (the thing that is actually percolating through the system) and then it becomes an “Active” one. The condition for a Waiting cell to transform into an Active one in the next step is that one or more of its “nearest cell neighbors” should be Active. An “Active” cell which is holding the Action Potential at a particular instant will be transformed into an “Inactive” one in the next time step. For a heart system this is consistent with the fact that there is a certain refractory period of the cells.
@@
The general rule of transformation for a cell may be described as: Waiting $\rightarrow$ Active $\rightarrow$ Inactive, or, using the numbers as $0\rightarrow 1\rightarrow 2$. However, in some specific cases an “Inactive” cell may become a “Waiting” one, i.e., $2\rightarrow 0$ is also allowed. This is similar to the tree regrowth at several sites in Forest Fire Model as discussed in [10, 11].
@@
  • $p_{act}$(inhibitory) by which an Waiting cell can become an Active one, i.e., $0\rightarrow 1$.
@@
The system is studied from two aspects. First, the number of arrival $N_{A}$ at the other end of the 2D matrix is observed with variation of the parameters $p_{act}$ and $p_{switch}$ in subsection A of the Results. If there is no Inactive cell, the AP reach all the cells of the final row in time $L$. However, when there is a distribution of 00 and $2$ cells, all the cells of the final row do not contain the AP at the same time but $1$s arrive at different cells at different times. At the final row the number of cells containing the AP is counted at each time. Second, the percolation model of the system is analyzed. The delay in arrival is measured and some interesting properties related to that are studied. In the same context the cluster distribution and several exponents related to the system are also studied. The corresponding results are shown in subsections B and C of the Results.
@@ Results
The Action Potential as initiated from one end of the 2D lattice of size $L$ percolates through the system. The number of arrivals at the other end, i.e., the active sites at that ends $N_{A}$ is checked with time $t$ We will study the number of arrivals $N_{A}$ with two conditions:
@@
Figure 1: $N_{A}$ vs $t$ with $L$ = 200, $p=0.70$, $p_{act}=1.0$ and $p_{switch}$ varying from 00 to $0.90$. In the inset $p_{switch}=0.90$ has been shown with equation 3 where $(N_{A})_{sat}\simeq 54$, $A=55$ and $B=0.05$.
The oscillatory behaviour can be explained in the following way. It is well known that when a fluid is flowing through a porous medium, there is some kind of relative motion [25] which generates a fluid shear [19], can be considered as a friction/damping. In the same way, when an electrical signal, here the AP, is moving through the heart, there must be some retardation present. For high $p_{switch}$, the motion is similar to an underdamped oscillatory one. This is obvious as for high $p_{switch}$, the signal more frequently finds a path to percolate, be it in forward or backward directions or sideways. Therefore number of arrival $N_{A}$ is never zero. However, for $p_{switch}=0.0$, it has been observed that the behaviour of $N_{A}$ is Gaussian. This signifies that for low $p_{switch}$ the number of active cells die down almost completely in time $L<t<2L$. Therefore for further AP signals there is no chance to pass down through the system. However, for higher $p_{switch}$, this is not true and even after a very long time there is always a chance for the next APs to pass through the system. This is because here we have $(N_{A})_{sat}>0$ always that indicateds a certain finite number of Active cells will always be present in the system. Even if we study the system for $t\rightarrow\infty$ the system still shows non-zero active sites.
@@
$\frac{d^{2}x}{dt^{2}}+2B\frac{dx}{dt}+\omega_{0}^{2}x=0$ (1)
where the coefficient $2B$ is related to damping and $\omega_{0}$ is the natural frequency of oscillation. For underdamped/oscillatory motion $B^{2}<\omega_{0}^{2}$ and the corresponding solution can be written as
$x(t)=A\cos(\omega t)\exp(-Bt)$ (2)
@@
$N_{A}(t)=(N_{A})_{sat}+A\cos(\omega{t^{\prime}}^{\nu})\exp(-Bt^{\prime})$ (3)
@@
We now fix $p_{switch}=0$ and study the system behaviour by varying $p_{act}$. It is already clear from Fig. 1 that the behaviour will be approximately Gaussian. In Fig. 2 the behaviour is shown for $0.85\leqslant p_{act}\leqslant 1.0$. The parameters related to the Gaussian behaviour is shown in Table 2 for $L=200$ by fitting with the following function:
$N_{A}(t)=(N_{A})_{max}e^{-\frac{1}{2}\left(\frac{t-\mu}{\sigma_{sd}}\right)^{2}}.$ (4)
@@
Figure 2: $N_{A}$ vs $t$ with $L=200$, $p_{switch}=1.0$ and $p_{act}$ varies from $0.85$ to $1.00$.
@@
Figure 3: Variations of the Gaussian parameters with $p_{act}$ where $L=50,100$ and $200$, (a) $\frac{(N_{A})_{max}}{L^{\eta_{A}}}$ versus $p_{act}$ with $\eta_{A}=0.7$, $m=1.16\pm 0.05$ and $c=-0.98\pm 0.05$, (b) $\frac{\mu}{L^{\eta_{\mu}}}$ versus $p_{act}$ with $\eta_{\mu}=1.0$, $a_{\mu}=1.26\pm 0.03$ and $b_{\mu}=-2.02\pm 0.20$ and (c) $\frac{\sigma}{L^{\eta_{\sigma}}}$ vs $p_{act}$ with $\eta_{\sigma}=1.0$, $a_{\sigma}=0.96\pm 0.03$ and $b_{\mu}=-7.82\pm 0.36$.
@@
Next we study the probability $P_{N}$ in terms of the fraction of percolating paths out of a $N_{T}$ number of trials against $p$. This is shown in Fig. 4 for $p_{act}=1.0,p_{switch}=0.0$ for different $L$. Here we have taken $N_{T}=1000$. In the inset of Fig. 4 the data collapse has been shown. It has been observed that the percolation threshold is at $p_{c}\simeq 0.59$ which is in good agreement with [26, 27, 28, 29]. The function is checked to be fitted with the function $\frac{(1+\tanh{\kappa x})}{2}$ where $\kappa=1.4$ and $x=\frac{p-p_{c}}{p_{c}}$. The same has also been studied for other combinations of $p_{act}$ and $p_{switch}$. As we decrease the value of $p_{act}$, the percolation threshold $p_{c}$ shifts towards higher side as the Waiting sites are becoming Active with less probability. If keeping $p_{act}=1.0$ we increase the value of $p_{switch}$ even slightly, $p_{c}$ shifts towards lower $p$. This is also in agreement with Fig. 1. These are shown in Fig.5 (a) for $p_{act}=0.80$, $p_{switch}=0.00$ and (b) $p_{act}=1.0$, $p_{switch}=0.03$. From Fig. 5(a) $p_{c}$ is found to be $0.72$ and from 5(b) it is found to be close to $0.12$. It is to be noted that in Fig. 5(b) there is no crossing and even for very small $p_{switch}$, $p_{c}$ drops rapidly.
Figure 4: $P_{N}$ vs $p$ with $p_{act}=1.0$, $p_{switch}=0.00$ where $L=20,30,50,70$ and $100$ in the inset data collapse has been shown as $\frac{(1+\tanh(\kappa x))}{2}$ with $n=1.5$, and $\kappa=1.4$
Figure 5: $P_{N}$ vs $p$ where $L=20,30,50,70$ and $100$ with $n=1.5$, $(a)$ $p_{act}=0.80$ and $p_{switch}=0.00$ in the inset data colaps has been shown as $\frac{(1+\tanh(\kappa x))}{2}$, $n=1.5$ and $\kappa=1.5$ $(b)$ $p_{act}=1.0$ and $p_{switch}=0.03$ in the inset data collapse has been shown as $\frac{(1+\tanh(\kappa(x-\epsilon))}{2}$, $n=3.5$, $\kappa=0.47$ and $\epsilon=6.31$.
@@
$\tau=\frac{\braket{l_{0}}}{L}=\frac{x_{0}}{L}$ (5)
@@
Of course, $\tau$ depends on percolation probability $p$, since below critical percolation $p_{c}$ signal cannot be transmitted through the system. The variation of $x_{0}$ as a function of $p$ is shown in Fig. 6 and $\tau$ as a function of $p$ is shown in the inset of the same Figure in log scale for $L=20,30,50,70$ and $100$.It has been observed that
$\tau\propto(p-p_{c})^{-u}$ (6)
The exponent $u$ is almost independent of $L$ and the approximated value is $u\simeq 0.20$. It is to be mentioned here that value of this exponent remains unaltered with variation of $p_{act}$.
Figure 6: $x_{0}$ vs $p$ with $p_{act}=1.00$ and $p_{switch}=0.0$ and $L=20,30,50,70$ and $100$ n the inset data collapse has been shown as $\tau$ vs $p$ with $u\simeq 0.20$.
@@
Now we would move on to find out some other relevant exponents for the system. It is well known that the cluster distribution for site percolation shows power law behaviour for $p=p_{c}$ as $n(s)_{p_{c}}\propto s^{-{\Delta}}$ where $\Delta$ is the Fisher exponent [11].
Figure 7: $n(s)$ vs $s$ with $L=500$, $p_{act}=1.00$ and $p_{switch}=0.0$ here $\Delta\simeq 1.96$.
Figure 8: $n(s)$ vs $s$ with $L=500$, $p_{act}=0.90$ and $p_{switch}=0.0$ here $\Delta\simeq 1.91$.
@@
Figure 9: $M$ vs $L$ with $p_{act}=1.00$ and $p_{switch}=0.0$ here $d_{f}=1.81\pm 0.01$.
Figure 10: Study of critical exponents, (a) $P_{\infty}L^{\xi}$ vs $p-p_{c}$ with $\beta=0.39\pm 0.04$ and $\xi=1.95$, (b) $\braket{s}$ vs $p-p_{c}$ with $\gamma=0.60\pm 0.02$ and (c) $\frac{s_{max}}{L}$ vs $|p-p_{c}|$ with $\sigma=1.00\pm 0.33$, which clearly satisfy the relation $\beta+\gamma=\frac{1}{\sigma}$.
Mass of the largest cluster $M$ for $p=p_{c}$ shows a typical behaviours $M\sim L^{d_{f}}$ where $L$ is the system size and $d_{f}$ is the fractal dimension. This is shown in Fig. 9. Here $d_{f}$ is found to be $d_{f}=1.81\pm 0.01$.
@@
The strength of the infinite cluster $P_{\infty}$ and the divergence of the mean cluster size $\langle s\rangle$ are another two important quantities. Those result in two more exponents $\beta$ and $\gamma$ which for our case happens to be $0.39\pm 0.04$ and $0.60\pm 0.02$ and are respectively shown in Fig. 10 (a) and (b).
From Fig. 7 another quantity is studied very often which is the cut-off cluster size $s_{max}$. This can be found from Fig. 7 (c) by using the well known form of $n_{s}\sim s^{-\Delta}\exp(-s/s_{max})$ for large $s$. The corresponding variation of cut-off cluster size $s_{max}$ for our system is shown in Fig. 10. From this we found the exponent $\sigma$ as $1.00\pm 0.33$, the relation being $s_{max}\propto(p-p_{c})^{-1/\sigma}$.
The value of the exponent $\alpha$ can be found out easily as $\sum_{0}^{\infty}n_{s}=(p-p_{c})^{2-\alpha}$. The value of $\alpha$ obtained in our case is close to $2.00$.
It is clear from the above facts that the Rushbrooke’s inequality $\alpha+2\beta+\gamma\geq 2$ is satisfied. Also the relation $\beta+\gamma=\frac{1}{\sigma}$ is verified here to be true.
@@
In the other part of the work we have studied the percolation model of the system. It has been observed that the critical probability for percolation $p_{c}$ is close to $0.5923$ for the $p_{switch}=0$ and $p_{act}=1$. The value of $p_{c}$ shifts towards higher and higher side as $p_{act}$ is made smaller and smaller. Keeping $p_{act}$ fixed at $1$ and changing $p_{switch}$ a little has severe effect on $p_{c}$ and it shifts abruptly towards 00. The tortuosity is studied and the related exponent is obtained as close to $0.2$. We also focused on the study of the exponents relevant to percolation theory. In this part we considered the parameters $p_{switch}=0$ and $p_{act}=1$ only. The cluster distribution is studied here as shown in Fig. 7. From this we checked the exponents $\alpha,\beta,\gamma,\sigma$ etc. The Rushbrooke Inequality $\alpha+2\beta+\gamma\geq 2$ and the relation $\beta+\gamma=1/\sigma$ has been observed to be satisfied.

### v2

@@ Abstract
We study the signal percolation through heart-like biological system. Starting from an initial distribution of waiting and inactive cells with probabilities $p$ and $(1-p)$ respectively, the signal propagation is observed in terms of active cells. As the signal enters the system from one end, the number of arrival of active sites at the other end is studied and analysis of the system behaviour is made by varying a few important parameters of the system like $p_{switch}$ (switching probability from inactive to waiting) and $p_{act}$ (switching probability from waiting to active). In this connection, the non-regular heart rhythms are discussed. Fraction of paths percolating through the system shows a transition from 00 to $1$ near $p=p_{c}$. Some other important quantities like tortuosity and cluster distribution are discussed. Several critical exponents have been obtained and compared the exponents of standard percolation.
@@ Introduction
Percolation is one of the simplest and fundamental models in statistical mechanics which is not always an exactly solvable model. In spite of its simple update rules, it can be applied to describe a wide variety of systems, e.g., fluid flow through porous media [1], electrical movement in neurons, fibrosis in organs, damage spreading in solids, disease spreading in a community, opinion formation in society or even for knowledge percolation [2, 3, 4, 5, 6, 7, 8, 9]. A very popular model of percolation is the forest fire model (FFM) introduced in [10, 11, 12] mainly as a 2D lattice studied as “directed percolation” (DP) [13, 14], isotropic percolation (IP) [11] and semi directed percolation (SDP) [15, 16, 17] in several variations.
@@
Several biological models have been used to study the heart system with the help of computation. Computational models in cardiology showed enormous scope in revealing diagnostic information related to cardiac disease. At the same time there are provisions that these models help the diagnosis to be cost-effective and to have low risk [5]. The cells can be activated by an external stimulus which initiates an action potential (AP). The cells after activation will go through a certain refractory period. The exact physiological processes are described in [22] in a detailed manner. In [6], the model of heart is studied with three states of cells: active, inactive and waiting in connection with the AP.
One of the simplified system to study a percolation problem is a $L\times L$ lattice, where there is a parameter called percolation probability $p$ with which a site could be open, i.e. can percolate “something”. This in turn indicates that a site is closed with $1-p$ probability, means unable to percolate. It is well known from different studies that below a critical percolation probability $p_{c}$ the system is unable to percolate, whereas above $p_{c}$ infinite percolation cluster can be formed [11, 23, 24, 25, 26, 27]. In connection to percolation there are a number of exponents which can be studied as well. Those exponents actually describes the system near criticality [11, 28, 29].
The heart-like system is studied here in the light of percolation theory. In this study, we tried to discuss several features related to the AP movement as in IP through the heart-like system by varying a few parameters in connection to the heart cells. Several critical exponents for the percolation model of the heart-like system are obtained and compared to the corresponding exponents obtained from standard percolation.
@@
We study a 2D square grid of $L\times L$ cells where initially the cells may be in two states, “waiting” and “inactive” as described in [6] also. A “waiting” cell is an open cell which is ready to hold anything percolating through the system. An “inactive” cell is however equivalent to a closed cell which is unable to percolate. We choose the system with an initial distribution of “waiting” cells with probability $p$ and “inactive” cells with probability $1-p$. A “waiting” cell can be activated with an AP (the “thing” that is actually percolating through the system) and then it becomes an “active” one. The condition for a waiting cell to transform into an active one in the next step is that one or more of its “nearest cell neighbors” should be active. An “active” cell which is holding the AP at a particular instant will be transformed into an “inactive” one in the next time step. For a heart system this is consistent with the fact that there is a certain refractory period of the cells.
@@
The general rule of transformation for a cell may be described as: waiting $\rightarrow$ active $\rightarrow$ inactive, or, using the numbers as $0\rightarrow 1\rightarrow 2$. However, in some specific cases an “inactive” cell may become a “waiting” one, i.e., $2\rightarrow 0$ is also allowed. This is similar to the tree regrowth at several sites in FFM as discussed in [10, 11].
@@
  • $p_{act}$(inhibitory) by which an waiting cell can become an active one, i.e., $0\rightarrow 1$,
@@
The system is studied from two aspects. First, in subsection A of the Results, the number of arrival $N_{A}$ at the other end of the 2D matrix is observed with variation of the parameters $p_{act}$ and $p_{switch}$. If there is no inactive cell, the AP reach all the cells of the final row in time $L$. However, when there is a distribution of 00 and $2$ cells, all the cells of the final row do not receive the AP at the same time but $1$s arrive at different cells at different times. At the final row the number of cells containing the AP is counted at each time. Second, in subsection B and C of the Results, the percolation model of the system is analyzed. The delay in arrival is measured and some interesting properties related to that are studied. In the same context the cluster distribution and several critical exponents related to the system are also studied.
@@ Results
The AP, as initiated from one end of the 2D lattice of size $L$, percolates through the system. The number of arrivals at the other end, i.e., the number of active sites $N_{A}$ at that end is checked with time $t$ We study the number of arrivals $N_{A}$ with two conditions:
@@
Figure 1: $N_{A}$ versus $t$ with $L$ = 200, $p=0.70$, $p_{act}=1.0$ and $p_{switch}$ varying from 00 to $0.90$. In the inset $p_{switch}=0.90$ has been shown with equation 3 where $(N_{A})_{sat}\simeq 54$, $A=55$ and $B=0.05$.
The oscillatory behaviour can be explained in the following way. It is well known that when a fluid is flowing through a porous medium, there is some kind of relative motion [32] which generates a fluid shear [19], can be considered as a friction/damping. In the same way, when an electrical signal, here the AP, is moving through the heart, there must be some retardation present. For high $p_{switch}$, the motion is similar to an underdamped oscillator. This is obvious as for high $p_{switch}$, the signal more frequently finds a path to percolate, be it in forward or backward directions or sideways. Therefore number of arrival $N_{A}$ is never zero. However, for $p_{switch}=0.0$, it has been observed that the behaviour of $N_{A}$ is Gaussian. This signifies that for low $p_{switch}$ the number of active cells die down almost completely in time $L<t<2L$. Therefore, there is no chance for further AP signals to pass down through the system. However, for higher $p_{switch}$, this is not true and even after a very long time, there is always a chance for the next APs to pass through the system. This is because here we have $(N_{A})_{sat}>0$ always. It indicates that, a certain finite number of active cells will always be present in the system. Even if we study the system for $t\rightarrow\infty$, the system still shows non-zero active sites.
@@
$\frac{d^{2}x}{dt^{2}}+2B\frac{dx}{dt}+\omega_{0}^{2}x=0,$ (1)
where the coefficient $2B$ is related to damping and $\omega_{0}$ is the natural frequency of oscillation. For underdamped/oscillatory motion $B^{2}<\omega_{0}^{2}$ and the corresponding solution can be written as,
$x(t)=A\cos(\omega t)\exp(-Bt),$ (2)
@@
$N_{A}(t)=(N_{A})_{sat}+A\cos(\omega{t^{\prime}}^{\nu})\exp(-Bt^{\prime}),$ (3)
@@
We now fix $p_{switch}=0$ and study the system behaviour by varying $p_{act}$. It is clear from Fig. 1 that the behaviour will be approximately Gaussian. In Fig. 2 the behaviour is shown for $0.85\leqslant p_{act}\leqslant 1.0$. The parameters related to the Gaussian behaviour is shown in Table 2 for $L=200$ by fitting with the following function:
$N_{A}(t)=(N_{A})_{max}e^{-\frac{1}{2}\left(\frac{t-\mu}{\sigma_{sd}}\right)^{2}},$ (4)
@@
Figure 2: $N_{A}$ versus $t$ with $L=200$, $p_{switch}=0.0$ and $p_{act}$ varies from $0.85$ to $1.00$.
@@
Figure 3: Variations of the Gaussian parameters with $p_{act}$ where $L=50,100$ and $200$, (a) $\frac{(N_{A})_{max}}{L^{\eta_{A}}}$ versus $p_{act}$ with $\eta_{A}=0.7$. For the fitted straight line, slope $m=1.161\pm 0.055$ and intercept $c=-0.976\pm 0.051$, (b) $\frac{\mu}{L^{\eta_{\mu}}}$ versus $p_{act}$ (in log scale) and the fitted function is $a_{\mu}p_{act}^{b_{\mu}}$ with $\eta_{\mu}=1.0$, $a_{\mu}=1.261\pm 0.026$ and $b_{\mu}=-2.023\pm 0.201$ and (c) $\frac{\sigma}{L^{\eta_{\sigma}}}$ versus $p_{act}$ (in log scale) and the fitted function is $a_{\sigma}p_{act}^{b_{\sigma}}$ with $\eta_{\sigma}=1.0$, $a_{\sigma}=0.960\pm 0.028$ and $b_{\sigma}=-7.819\pm 0.358$.
@@
Next we study the fraction of percolating paths out of $N_{T}$ number of trials, termed as probability $P_{N}$ here, as a function of $p$. This is shown in Fig. 4 for $p_{act}=1.0,p_{switch}=0.0$ for different $L$. Here we have taken $N_{T}=10^{5}$. In the inset of Fig. 4 the data collapse has been shown. It has been observed that the percolation threshold is at $p_{c}=0.5925\pm 0.0001$ which is in good agreement with [23, 24, 26, 27]. The function is checked to be fitted with the function $\frac{(1+\tanh{\kappa x})}{2}$ where $\kappa=0.96$, $x=(\frac{p-p_{c}}{p_{c}})L^{1/{\nu}}$ and $\nu=\frac{4}{3}$. The same has also been studied for other combinations of $p_{act}$ and $p_{switch}$. As we decrease the value of $p_{act}$, the percolation threshold $p_{c}$ shifts towards higher side, as the waiting sites are becoming active with less probability. If keeping $p_{act}=1.0$ we increase the value of $p_{switch}$ even slightly, $p_{c}$ shifts towards lower $p$. This is also in agreement with Fig. 1. These are shown in Fig.5 (a) for $p_{act}=0.80$, $p_{switch}=0.00$ and (b) $p_{act}=1.0$, $p_{switch}=0.03$. From Fig. 5(a) $p_{c}$ is found to be $0.7157$ and from 5(b) it is found to be close to $0.12$. It is to be noted that in Fig. 5(b) there is no crossing and even for very small $p_{switch}$, $p_{c}$ drops rapidly.
Figure 4: $P_{N}$ versus $p$ with $p_{act}=1.0$, $p_{switch}=0.00$ where $L=20,30,50,70,100$ and $200$ in the inset data collapse has been shown as $\frac{(1+\tanh(\kappa x))}{2}$ with $\nu=\frac{4}{3}$ and $\kappa=0.96$.
Figure 5: $P_{N}$ versus $p$ where $L=20,30,50,70,100$ and $200$, $(a)$ $p_{act}=0.80$ and $p_{switch}=0.00$. In the inset data collapse has been shown and the fitted function is $\frac{(1+\tanh(\kappa x))}{2}$, $\nu=\frac{4}{3}$ and $\kappa=1.0$ $(b)$ $p_{act}=1.0$ and $p_{switch}=0.03$. In the inset data collapse has been shown and the fitted function is $\frac{(1+\tanh(\kappa(x-\epsilon))}{2}$, $\nu=3.7$, $\kappa=0.52$ and $\epsilon=5.89$.
@@
$\tau=\frac{\braket{l_{0}}}{L}=\frac{x_{0}}{L},$ (5)
@@
Of course, $\tau$ depends on percolation probability $p$, since below critical percolation $p_{c}$ signal cannot be transmitted through the system. The variation of $x_{0}$ as a function of $p$ is shown in Fig. 6 and $\tau$ as a function of $p$ is shown in the inset of the same Figure in log scale for $L=20,30,50,70$ and $100$. It has been observed that,
$\tau\propto(p-p_{c})^{-u}.$ (6)
The exponent $u$ is almost independent of $L$ and the approximated value is $u\simeq 0.205\pm 0.003$. It is to be mentioned here that value of this exponent remains unaltered with variation of $p_{act}$.
Figure 6: $x_{0}$ versus $p$ with $p_{act}=1.00$ and $p_{switch}=0.0$ and $L=20,30,50,70$ and $100$ n the inset data collapse has been shown as $\tau$ versus $p$ with $u=0.205\pm 0.003$.
@@
Now we would move on to find out some other relevant exponents for the system. It is well known that, the cluster distribution for site percolation shows power law behaviour for $p=p_{c}$ as $n(s)_{p_{c}}\propto s^{-{\Delta}}$, where $\Delta$ is the Fisher exponent [11].
Figure 7: $n(s)$ versus $s$ with $L=500$, $p_{act}=1.00$ and $p_{switch}=0.0$ here $\Delta=1.964\pm 0.030$.
Figure 8: $n(s)$ versus $s$ with $L=500$, $p_{act}=0.90$ and $p_{switch}=0.0$ here $\Delta=1.911\pm 0.002$.
@@
Figure 9: $M$ versus $L$ with $p_{act}=1.00$ and $p_{switch}=0.0$ here $d_{f}=1.811\pm 0.011$.
Figure 10: Study of critical exponents : (a) $P_{\infty}L^{\xi}$ versus $p-p_{c}$ with $\bar{\beta}=0.392\pm 0.045$ and $\xi=1.85$, which gives $\beta=2.859$, (b) $\braket{s}$ versus $p-p_{c}$ with $\gamma=0.600\pm 0.021$ and (c) $\frac{s_{max}}{L}$ versus $|p-p_{c}|$ with $\bar{\sigma}=0.757\pm 0.150$, which gives $\sigma=0.377$. These exponents approximately satisfy the relation $\beta+\gamma=\frac{1}{\sigma}$.
Mass of the largest cluster $M$ for $p=p_{c}$ shows a typical behaviours $M\sim L^{d_{f}}$ where $L$ is the system size and $d_{f}$ is the fractal dimension. This is shown in Fig. 9. Here $d_{f}$ is found to be $d_{f}=1.811\pm 0.011$.
@@
The strength of the infinite cluster $P_{\infty}$ and the divergence of the mean cluster size $\langle s\rangle$ are another two important quantities. Those result in two more exponents $\beta$ and $\gamma$ which for our case happens to be $2.859$ and $0.600\pm 0.021$ and are respectively shown in Fig. 10 (a) and (b).
From Fig. 7 another quantity is studied very often which is the cut-off cluster size $s_{max}$. This can be found from Fig. 7 (c) by using the well known form of $n_{s}\sim s^{-\Delta}\exp(-s/s_{max})$ for large $s$. The corresponding variation of cut-off cluster size $s_{max}$ for our system is shown in Fig. 10. From this we found the exponent $\sigma$ as $0.377$, the relation being $s_{max}\propto(p-p_{c})^{-1/\sigma}$.
The value of the exponent $\alpha$ can be found out easily as $\sum_{0}^{\infty}n_{s}=(p-p_{c})^{2-\alpha}$. The value of $\alpha$ obtained in our case is close to $2.0004$.
It is clear from the above facts that the Rushbrooke’s inequality $\alpha+2\beta+\gamma\geq 2$ is satisfied. Also the relation $\beta+\gamma=\frac{1}{\sigma}$ is verified here to be approximately true.
@@
In the next part of the work, we have studied the percolation model of the system and analyzed the critical exponents for the same. It has been observed that the critical probability for percolation $p_{c}$ is close to $0.5925$ for the $p_{switch}=0$ and $p_{act}=1$.
$p_{act}$ $p_{switch}$ $p_{c}$
1.00 0.00 0.5925
0.95 0.00 0.6185
0.90 0.00 0.6476
0.80 0.00 0.7157
1.00 0.03 0.12
Table 3: Approximate values of $p_{c}$ for different $p_{act}$ and $p_{switch}$
The value of $p_{c}$ shifts towards higher and higher side as $p_{act}$ is made smaller and smaller. Keeping $p_{act}$ fixed at $1$ and changing $p_{switch}$ a little has severe effect on $p_{c}$ and it shifts abruptly towards 00. The way $p_{c}$ is changing is shown in Table 3.
Model $\Delta$ $\alpha$ $\beta$ $\gamma$ $\sigma$ $\nu$
Heart Model
$p_{act}=1.0$ $1.964$ $2.0004$ $2.859$ $0.600$ $0.377$ $4/3$
$p_{switch}=0.0$
Standard Model $187/91$ $-2/3$ $5/36$ $43/18$ $36/81$ $4/3$
Table 4: Comparison of critical exponents.
The tortuosity is studied and the related exponent $u\simeq 0.205$ is obtained. We also studied the exponents relevant to percolation theory. In this part, we considered the parameters $p_{switch}=0$ and $p_{act}=1$ only. The cluster distribution is studied here as shown in Fig. 7. From this, we checked the exponents $\alpha,\beta,\gamma,\sigma$ etc. which are shown in Table 4. It has been observed that a few exponents are quite close to those obtained in case of standard percolation. However, other exponent values deviate from the standard case. This may be due to the transformation of cells as $0\rightarrow 1\rightarrow 2\rightarrow 0$, which is a system specific property and therefore is not exactly similar to the standard percolation. The Rushbrooke Inequality $\alpha+2\beta+\gamma\geq 2$ and the relation $\beta+\gamma=1/\sigma$ has been observed to be closely satisfied.